# Improving Black-box Robustness with In-Context Rewriting

### v2

@@ Abstract
Machine learning models often excel on in-distribution (ID) data but struggle with unseen out-of-distribution (OOD) inputs. Most techniques for improving OOD robustness are not applicable to settings where the model is effectively a black box, such as when the weights are frozen, retraining is costly, or the model is leveraged via an API. Test-time augmentation (TTA) is a simple post-hoc technique for improving robustness that sidesteps black-box constraints by aggregating predictions across multiple augmentations of the test input. TTA has seen limited use in NLP due to the challenge of generating effective natural language augmentations. In this work, we propose LLM-TTA, which uses LLM-generated augmentations as TTA’s augmentation function. LLM-TTA outperforms conventional augmentation functions across sentiment, toxicity, and news classification tasks for BERT and T5 models, with BERT’s OOD robustness improving by an average of 4.30 percentage points without regressing average ID performance. We explore selectively augmenting inputs based on prediction entropy to reduce the rate of expensive LLM augmentations, allowing us to maintain performance gains while reducing the average number of generated augmentations by 57.76%. LLM-TTA is agnostic to the task model architecture, does not require OOD labels, and is effective across low and high-resource settings. We share our data111https://huggingface.co/datasets/Kyle1668/LLM-TTA-Augmentation-Logs, models222https://huggingface.co/collections/Kyle1668/, and code333https://github.com/Kyle1668/LLM-TTA for reproducibility.
@@ 1 Introduction
LLM-TTA Improves OOD Robustness. ICR improves a BERT classifier’s absolute accuracy on OOD data by an average of 4.86% for sentiment, 6.85% for toxicity, and 1.18% for news topics, all with minimal regression to ID performance. For toxicity, BERT’s ID accuracy improves by 2.99%, suggesting that LLM-TTA can improve both ID and OOD performance in some settings. Ablating training set size (Section 5.4) shows that LLM-TTA is useful in data-scarce and rich settings.
@@
TTA with Conventional Augmentation Functions Often Hurts Performance. In contrast to LLM-TTA, TTA with conventional augmentation functions generally hurt both ID and OOD performance. Back-translation is the best-performing conventional augmentation functions, with BERT’s OOD robustness improving by an average of 2.85 percentage points while word insertion regresses performance by –0.22 points and substitution by –0.17 points averaged across tasks.
@@
Selectively Augmenting High-Entropy Test Inputs Improves Efficiency. We can reduce the rate of expensive LLM augmentations by only augmenting test inputs in which the task model is uncertain in its prediction. For BERT as the task model and ICR as the augmentation function, we reduce the percentage of test inputs requiring augmentation by an average of 57.76% while still improving robustness. The uncertainty threshold is only determined from ID statistics.
@@ 3 LLM-TTA: Generating Faithful Augmentations With LLMs
3.1 Augmenting Out-of-Distribution Inputs
@@
3.2 Rewriting OOD Inputs with Language Models
@@ 4.1 Datasets
We consider three sentiment classification datasets from the BOSS benchmark (Yuan et al., 2023). Each is a three-way classification task where the labels are positive, neutral, and negative. The ID dataset consists of Amazon reviews (McAuley & Leskovec, 2013), while the three OOD datasets are DynaSent (Potts et al., 2020), SST-5 (Socher et al., 2013), and SemEval (Nakov et al., 2016). Yuan et al. (2023) selected these three OOD shifts since their centroids had low cosine similarity with the ID (Amazon) centroid.
@@
We also leverage the toxicity task in the BOSS benchmark. Toxicity detection is framed as a binary classification task between non-toxic (negative) and toxic (positive). The ID dataset is Civil Comments (Borkan et al., 2019), a collection of users’ comments on articles on the Civil News platform. The OOD datasets are AdvCivil, an adversarial version of Civil Comments introduced in the benchmark, as well as existing datasets ToxiGen (Hartvigsen et al., 2022) and Implicit Hate (Elsherief et al., 2021). These shifts were similarly selected by Yuan et al. (2023) since they have low cosine similarity with the ID centroid.
@@
AG News (Zhang et al., 2015) is a four-class news topic classification problem where the model is tasked with determining whether a given news article pertains to “World," “Sports," “Business," or “Sci/Tech." This task has a single OOD dataset, a novel dataset that is composed of the AG News test set style-transferred to resemble social media posts. Since each entry in the ID evaluation set has a corresponding style transferred entry in the OOD set, we can isolate the effect that differences in writing style have on performance. Additional details on how this dataset was created can be found in Appendix A.8.
@@ 4.2 Task Models
We study black-box robustness for three task models. First, we finetune BERT (Devlin et al., 2019) and T5-Large (Raffel et al., 2020) for each ID dataset. Second, we consider Falcon-7b, an instruction-tuned LLM (Almazrouei et al., 2023). To make predictions with Falcon, we use 16-shot prompt with randomly selected ID exemplars with equal numbers per class. Finetuning details are in Appendix A.5, and Falcon prompting details are in Appendix A.10.
@@ 4.4 TTA Settings
We use Stable Beluga 2-7B (SB2) to generate augmentations. SB2 is a LLama 2 (Touvron et al., 2023) model fine-tuned with additional instruction tuning. Figure 3 details the prompts for the LLM-TTA paraphrasing and ICR methods. These prompts and the test input are passed to the model, resulting in four stochastic augmentations. Specifically, we use the HuggingFace generation API with four return sequences, four beam groups, four beams, and a diversity penalty of 0.5. The maximum number of new tokens is dynamically set to be four times the number of tokens in the test input. ICR uses 16 randomly selected unlabeled exemplars balanced across classes sourced from the ID training set.
@@ 5.1 LLM-TTA Improves OOD Robustness
Generating augmentations with LLMs outperforms all other augmentation functions for BERT and T5 across each OOD task, as shown in Table 5.1. We report the average accuracy across the multiple OOD shifts for sentiment and toxicity. ICR is the best-performing augmentation function, with BERT’s OOD performance improving by an average of 4.30 percentage points and T5’s by 2.81 percentage points. Falcon benefits less with LLM-TTA regressing performance on sentiment and toxicity for an overall net regression of –3.01 percentage points.
@@
LLM-TTA can improve task model robustness without modifying the task model’s weights or training regime. TTA’s flexibility allows it to be slotted into existing systems. The degree to which TTA improves robustness is contingent on the augmentation function, with whole-text augmentation functions outperforming word-level augmentations. While we observe consistent improvements, more progress is needed to make models robust to OOD shifts as test time.
@@
| Substitute (Lu et al., 2022) | 51.78% | 55.07% | 44.27% | 53.37% | 57.65% | 58.78% | 88.86% | 89.13% | 24.88% |
| Translate (Sennrich et al., 2015) | 52.60% | 56.86% | 44.98% | 61.65% | 63.51% | 58.97% | 88.82% | 89.04% | 26.36% |
@@ 5.3 Selective Augmentation Improves Efficiency
Table 5.3 demonstrates that selective augmentation can preserve most performance gains while drastically reducing the number of expensive LLM calls. ICR tends to benefit more than paraphrasing. BERT’s robustness on sentiment is improved by 4.16 points when only augmenting 56.51% of inputs compared to 4.86 points when augmenting every input. Similarly, the majority of gains are still realized for toxicity while reducing the augmentation rate by roughly a third. News only preserved half of the performance gains yet reduced the augmentation rate by over 95% for ICR and paraphrasing.

### v3

@@ Abstract
Machine learning models for text classification often excel on in-distribution (ID) data but struggle with unseen out-of-distribution (OOD) inputs. Most techniques for improving OOD robustness are not applicable to settings where the model is effectively a black box, such as when the weights are frozen, retraining is costly, or the model is leveraged via an API. Test-time augmentation (TTA) is a simple post-hoc technique for improving robustness that sidesteps black-box constraints by aggregating predictions across multiple augmentations of the test input. TTA has seen limited use in NLP due to the challenge of generating effective natural language augmentations. In this work, we propose LLM-TTA, which uses LLM-generated augmentations as TTA’s augmentation function. LLM-TTA outperforms conventional augmentation functions across sentiment, toxicity, and news classification tasks for BERT and T5 models, with BERT’s OOD robustness improving by an average of 4.48 percentage points without regressing average ID performance. We explore selectively augmenting inputs based on prediction entropy to reduce the rate of expensive LLM augmentations, allowing us to maintain performance gains while reducing the average number of generated augmentations by 57.74%. LLM-TTA is agnostic to the task model architecture, does not require OOD labels, and is effective across low and high-resource settings. We share our data111https://huggingface.co/datasets/Kyle1668/LLM-TTA-Augmentation-Logs, models222https://huggingface.co/collections/Kyle1668/, and code333https://github.com/Kyle1668/LLM-TTA for reproducibility.
@@ 1 Introduction
Our study is composed of nine public datasets and one novel synthetic dataset across sentiment analysis, toxicity detection, and new topic classification. We consider the dataset used for model optimization as ID, while the other challenging evaluation sets are considered OOD. The experimental setup and its limitations are discussed in Sections 4.1 and 6.
@@
LLM-TTA Improves OOD Robustness. ICR improves a BERT classifier’s absolute accuracy on OOD data by an average of 5.12% for sentiment, 7.18% for toxicity, and 1.15% for news topics, all with minimal regression to ID performance. For toxicity, BERT’s ID accuracy improves by 2.99%, suggesting that LLM-TTA can improve both ID and OOD performance in some settings. Ablating training set size (Section 5.5) shows that LLM-TTA is useful in data-scarce and rich settings.
@@
TTA with Conventional Augmentation Functions Often Hurts Performance. In contrast to LLM-TTA, TTA with conventional augmentation functions generally hurt both ID and OOD performance. Back-translation is the best-performing conventional augmentation functions, with BERT’s OOD robustness improving by an average of 2.85 percentage points while word insertion regresses performance by –0.29 points and substitution by –0.53 points averaged across tasks.
@@
Selectively Augmenting High-Entropy Test Inputs Improves Efficiency. We can reduce the rate of expensive LLM augmentations by only augmenting test inputs in which the task model is uncertain in its prediction. For BERT as the task model and ICR as the augmentation function, we reduce the percentage of test inputs requiring augmentation by an average of 57.74% while still improving robustness. The uncertainty threshold is only determined from ID statistics.
@@ 3 LLM-TTA: Generating Faithful Augmentations With LLMs
3.1 Augmenting Test Inputs
@@
3.2 Rewriting Test Inputs with Language Models
@@ 4.1 Datasets
We refer to the dataset that we optimize the model for as ID. The challenging splits from other datasets, which the ID model struggles to generalize, are considered OOD. We do not investigate the specific properties that make these datasets challenging. This abstract notion of OOD contrasts with other works (Hendrycks et al., 2020a; Koh et al., 2020). See Section 6 for further discussion.
@@
We consider three sentiment classification datasets from the BOSS benchmark (Yuan et al., 2023). Each is a three-way classification task where the labels are positive, neutral, and negative. We keep this benchmark’s selection of ID and OOD splits. The ID dataset consists of Amazon reviews (McAuley & Leskovec, 2013), while the three OOD datasets are DynaSent (Potts et al., 2020), SST-5 (Socher et al., 2013), and SemEval (Nakov et al., 2016). Yuan et al. (2023) selected these three OOD shifts since their centroids had low cosine similarity with the ID (Amazon) centroid.
@@
We also leverage the toxicity task in the BOSS benchmark. Toxicity detection is framed as a binary classification task between non-toxic (negative) and toxic (positive). We similarly rely on the ID and OOD dataset selections from Yuan et al. (2023). The ID dataset is Civil Comments (Borkan et al., 2019), a collection of users’ comments on articles on the Civil News platform. The OOD datasets are AdvCivil, an adversarial version of Civil Comments introduced in the benchmark, as well as existing datasets ToxiGen (Hartvigsen et al., 2022) and Implicit Hate (Elsherief et al., 2021). As with sentiment, these shifts were selected by Yuan et al. (2023) since they have low cosine similarity with the ID centroid.
@@
AG News (Zhang et al., 2015) is a four-class news topic classification problem where the model is tasked with determining whether a given news article pertains to “World,” “Sports,” “Business,” or “Sci/Tech.” This task has a single OOD dataset, a novel dataset that is composed of the AG News test set style-transferred to resemble social media posts. Since each entry in the ID evaluation set has a corresponding style transferred entry in the OOD set, we can isolate the effect that differences in writing style have on performance. Additional details on how this dataset was created can be found in Appendix A.7.
@@ 4.2 Task Models
We study black-box robustness for three task models. First, we fine-tune BERT (Devlin et al., 2019) and T5-Large (Raffel et al., 2020) for each ID dataset. Second, we consider Falcon-7b, an instruction-tuned LLM (Almazrouei et al., 2023). To make predictions with Falcon, we use 16-shot prompt with randomly selected ID exemplars with equal numbers per class. Fine-tune details are in Appendix A.5, and Falcon prompting details are in Appendix A.9.
@@ 4.3 Baseline Augnmention Functions
4.4 Multiple Evaluation Runs
To ensure that performance is robust across generated augmentation, we run our OOD evals four times using various seeds: 3, 16, 46, and 58. The results reported for OOD shifts are the mean values derived from these four runs, as well as the standard deviation. Each augmentation function is non-deterministic, which results in differing augmentations being generated across runs. This approach allows us to assess how consistently each augmentation function affects performance. We conduct a single run for the ID datasets.
@@ 4.5 TTA Settings
We use Stable Beluga 2-7B (SB2) to generate augmentations. SB2 is a LLama 2 (Touvron et al., 2023) model fine-tuned with additional instruction tuning. Figure 3 details the prompts for the LLM-TTA paraphrasing and ICR methods. These prompts and the test input are passed to the model, resulting in four stochastic augmentations. ICR uses 16 randomly selected unlabeled exemplars balanced across classes sourced from the ID training set. We further describe the decoding details in Appendix A.9.
@@ 5.1 LLM-TTA Improves OOD Robustness
Generating augmentations with LLMs outperforms all other augmentation functions for BERT and T5 across each OOD task, as shown in Table 5.1. We report the average accuracy across the multiple OOD shifts for sentiment and toxicity which is then average across four experiment runs. ICR is the best-performing augmentation function, with BERT’s OOD performance improving by an average of 4.48 percentage points and T5’s by 2.66 percentage points. Falcon benefits less with LLM-TTA regressing performance on sentiment and toxicity for an overall net regression of –0.67 percentage points.
@@
LLM-TTA can improve task model robustness without modifying the task model’s weights or training regime. TTA’s flexibility allows it to be slotted into existing systems. The degree to which TTA improves robustness is contingent on the augmentation function, with whole-text augmentation functions outperforming word-level augmentations. The modest gains, even with a powerful augmentation function, suggest that augmentation quality may not be the only bottleneck for improving black-box robustness. While we observe improvements, further progress is needed to make models more robust to OOD shifts as test time.
@@
Substitute (Lu et al., 2022) | 51.44% | 54.15% | 43.56% | 52.70% | 56.99% | 58.87% | 88.77% | 89.08% | 24.78% |
Translate (Sennrich et al., 2015) | 52.41% | 56.60% | 44.02% | 62.11% | 63.40% | 59.08% | 88.53% | 89.13% | 26.88% |
@@ 5.2 LLM-TTA Can Improve ID Performance
5.3 LLM-TTA Can Outperform the LLM
LLM-TTA augments every test input by default. In practice, using the augmentation LLM directly for the task may be advantageous if it can outperform the task model. In this section, we evaluate Stable Beluga 2 (SB2), the LLM we use for augmentation in our experiments, directly on the tasks. SB2 is prompted with the same task templates and exemplars as Falcon detailed in Appendix A.9.
We report our results in Table 5.3. SB2 underperforms BERT on all ID datasets and most OOD sets. SB2 overall outperforms BERT and T5 on the OOD toxicity detection and outperforms Falcon across most tasks. These results show that LLM-TTA can still improve a task model’s robustness even when the LLM itself underperforms the task model. However, using the LLM directly may be practical in some settings.
Sentiment | Toxicity | News Tweets | ||||
Augmentation | ID | OOD | ID | OOD | ID | OOD |
LLM-TTA ICR: BERT | 90.85% | 57.17% | 91.45% | 61.06% | 93.53% | 89.72% |
LLM-TTA ICR: T5 | 90.78% | 59.75% | 91.87% | 63.83% | 91.99% | 90.30% |
LLM-TTA ICR: Falcon | 90.35% | 43.79% | 19.00% | 58.11% | 29.47% | 27.82% |
[…]
@@ 5.4 Selective Augmentation Improves Efficiency
Table 5.4 demonstrates that selective augmentation can preserve most performance gains while drastically reducing the number of expensive LLM calls. ICR tends to benefit more than paraphrasing. BERT’s robustness on sentiment is improved by 4.33 points when only augmenting 56.53% of inputs compared to 5.12 points when augmenting every input. Similarly, the majority of gains are still realized for toxicity while reducing the augmentation rate by roughly a third. News only preserved half of the performance gains yet reduced the augmentation rate by over 95% for ICR and paraphrasing.